OpenStudy (anonymous):

If 2x^2 + x = 6, then which of the following could be the value of 2x^2 – x?

OpenStudy (anonymous):

\[\LARGE 2x^2+x-6=0 \quad use \quad x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] since it's quadratic equation..

OpenStudy (anonymous):

a=2 b=1 c=-6

OpenStudy (anonymous):

\[\LARGE x_1=-2 \quad ,\quad x_2=\frac32\]

OpenStudy (anonymous):

now substitute the values into the second equation... and find out what the solutions are... (If that's what is required...)

OpenStudy (anonymous):

is it 4 ?

OpenStudy (anonymous):

if you substitute x=3/2 you'll get 3 if you substitute x=-2 you'll get...10

OpenStudy (anonymous):

\[\LARGE 2\cdot \left(\frac32\right)^2-\frac32\] continue... and do the same with the second one too.. ;)

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

I got it, great thanks! :D